### v1

@@ 4.1.2 Tokenization
One of the primary limitations of the current mainstream Byte Pair Encoding (BPE)[49] tokenization method lies in its handling of numerical values. BPE can introduce ambiguity and inconsistency when tokenizing numbers:
  • Ambiguity in Tokenized Numerical Values: The number 12345 could potentially be segmented into ’12’, ’34’, and ’5’, resulting in the loss of the original numeric value’s inherent meaning.
  • Inconsistency in Tokenization of Numerical Values: Since BPE is a statistical method, it may divide numbers into different sub-units based on frequency patterns. This can lead to inconsistencies where the same number is tokenized differently depending on the context. For instance, the number 12345 could be segmented into ”12”, ”34”, and ”5” in one context, but in another context, it might be broken down into ”1”, ”23”, and ”45”.
  • Discontinuity in Token IDs for Numerical Values: Moreover, the BPE (Byte Pair Encoding) tokenization process frequently leads to fragmented token IDs for numerical entities. For instance, the digits ’7’ and ’8’ could be fragmented into token IDs 4779 and 5014, respectively. Such fragmentation poses a challenge in the efficient management and processing of numerical data, especially in contexts where sequential or patterned token IDs are advantageous. The absence of a coherent and continuous correlation between numerical values and their corresponding token IDs can exacerbate complexity in subsequent tasks, including numerical analysis and the deployment of machine learning models.
Similarly, single-digit tokenization[17], while simple, also suffers from significant drawbacks:
  • Discontinuous Token IDs for Numerical Values: Single-digit tokenization treats each digit as a separate token, leading to discontinuous token IDs for multi-digit numbers. In the case of the number 12345, it would be broken down into separate tokens: ’1’, ’2’, ’3’, ’4’, and ’5’, each receiving its own token ID. This segmentation may cause the loss of numerical relationships among the digits, challenging the preservation of the original numeric value’s significance..
@@
  • Preservation of Numerical Meaning: By converting numbers into byte sequences, the inherent meaning of the original numeric value is preserved. Each number is represented consistently, regardless of context.
  • Consistent Tokenization: The byte tokenization method ensures that the same number is always tokenized in the same way, avoiding the inconsistencies seen in BPE.
  • Continuous Token IDs: Byte tokenization results in a continuous and systematic mapping of numerical values to token IDs.By employing specific computer encoding techniques, numbers like 7 and 8 can be converted into sequential token IDs, thereby streamlining the management and processing of numerical data. This encoding method transforms numerical values into byte sequences, ensuring a consistent mapping from numbers to their corresponding token IDs.
Our Byte Tokenization method operates on the principle of encoding all input data into byte arrays, which serve as a common ground for further processing. The general workflow consists of three main steps:
@@
    • Text: Characters are translated into their corresponding Unicode byte values. In the case of the string ’hello’, this translation results in the byte array $[104,101,108,108,111]$.
    • Numbers: Numbers can be processed in two distinct manners. First, a number can be treated as a string and subsequently encoded using UTF-8. For instance, the number 12345 would be represented as the string ”12345” and then encoded into a byte array[49, 50, 51, 52, 53]. This method is particularly useful for preserving the exact format and any leading zeros that might be significant. Second, a number can be converted into its numerical form (e.g., an integer) and then transformed into a byte array. The integer 12345 can be directly converted into a byte array representation[0, 0, 0, 0, 11, 17, 20, 21]. This approach is more efficient for arithmetic operations and when the numerical value itself is crucial.
@@
    • Image:Images are typically represented as pixel values, where each pixel can have one or more bytes depending on the color depth. For grayscale images, each pixel is represented by a single byte (8-bit), while for color images (e.g., RGB), each pixel is represented by three bytes (24-bit). The conversion process involves reading the image file and extracting the pixel values into a byte array. Consider a small 2x2 grayscale image; its pixel values, represented as $[128,64,192,255]$, would correspond to the byte array $[128,64,192,255]$ upon conversion.
@@ 4.1.3 Training
The tokenization process varies between Patch Version and Non-patch Version. Patch Version involves converting attribute values into byte sequences and then segmenting these sequences into byte patches. Non-patch Version does not require cutting the byte sequences into patches. Instead, each byte is treated individually, and the entire sequence is processed as a continuous stream of bytes.
As a result, the total number of byte patches generated depends on the Patch Size and the length of the original byte sequence.
@@
We can present a simplified Figure 7 to illustrate how a jet is converted into a series of byte tokens. A jet consists of several particles, each described by a set of attribute values. These attribute values of different particles are uniformly converted into byte sequences. The byte sequences are then segmented into byte patches of a fixed length defined by the Patch Size, resulting in a total of Patch Sequence Length byte patches. Each byte patch has an input dimension equal to the patch size, and these patches form the input bytes for the model.
Data preprocessing can be implemented in either Streaming or Non-streaming mode, depending on the dataset size and available memory. In Non-streaming mode, all data is loaded and processed at once, making this approach suitable for smaller datasets or when sufficient memory is available. The main steps are:
  1. 1.
    Splitting the text into individual byte patches.
  2. 2.
    Ensuring uniform sequence length by padding shorter sequences and truncating longer ones.
  3. 3.
    Saving the preprocessed data to files for later use in training.
Streaming mode processes data in batches as it is read from the source, which is ideal for large datasets or limited memory scenarios. This method optimizes memory usage and processing efficiency. Key steps in streaming data preprocessing include:
  1. 1.
    Reading data in batches, loading only a portion of it into memory at any given time.
  2. 2.
    Tokenizing text in real-time as it streams in, avoiding the need to load the entire dataset.
  3. 3.
    Performing padding and truncation operations in real-time to maintain uniform sequence length.
  4. 4.
    Saving the preprocessed data in batches to files or databases, preventing memory overflow.
Training Objective The JOI (Jet Origin Identification) task is ideally suited for classification models as it entails assigning input data to specific categories. This classification task requires the model to accurately identify and differentiate between various types of jets based on their features and attributes. Since the model need to determine what type of jet this is,the label are ”B-jet”, ”B-bar-jet”, ”C-jet”, ”C-bar-jet”, ”S-jet”, ”S-bar-jet”, ”U-jet”, ”U-bar-jet””D-jet”, ”D-bar-jet” and ”G-jet”.
@@ 4.2 ParticleNet and Particle Transformer
The core concept of ParticleNet [50] is the EdgeConv operation, which begins by representing a jet as a graph, where the vertices correspond to particles, and the edges are constructed by connecting each particle to its k nearest neighbors. The Multi-Layer Perceptron (MLP) is then applied to these edges three times, followed by updating the features of each vertex based on the information from its connected edges. This iterative process, through the EdgeConv operation, enables the exchange of information between vertices and edges. The ParticleNet architecture consists of three EdgeConv blocks, followed by a channel-wise global average pooling block, which aggregates information from all vertices in the graph. Two fully connected layers are then applied, with a softmax function to generate predictions for each jet.
The Particle Transformer [46] employs an encoder-only architecture, followed by a projection layer to output classification scores. The inputs consist of particle features for each individual particle and interaction terms for every pair of particles. Both particle and interaction inputs are passed through an MLP for embedding. Unlike Transformers in Natural Language Processing or Computer Vision, the Particle Transformer does not include positional encoding, as the particles in a jet are permutation invariant. The particle embeddings are processed through a stack of L particle attention blocks, where multi-head self-attention is applied to produce new embeddings. The interaction embeddings are used to augment the scaled dot-product attention, serving as a bias to the pre-softmax attention weights. After the encoding process, a class token—a randomly initialized vector—gathers information via attention to all particles processed by the final attention block. In other words, the particles, after being processed by the final attention block, are projected onto the class token. Finally, the class token is passed through a single-layer MLP, followed by a softmax layer, to generate the final classification scores.
For the JoI task, both ParticleNet and Particle Transformer conceptualize a jet as a graph, where each particle is represented as a node. In ParticleNet, node features are updated iteratively based on interactions with their k-nearest neighbors. This involves constructing edges to the nearest neighbors and aggregating the associated edge features. By applying this operation over multiple layers, each node incrementally incorporates information from the broader graph structure. In contrast, Particle Transformer employs an attention mechanism [51], enabling each node to interact directly with every other node. This approach effectively treats the graph as fully connected, where the attention scores can be interpreted as dynamic edge weights. Following this message-passing phase, during which information propagates across the graph, both models condense the features of all nodes into a single vector, albeit using distinct aggregation methods. This vector is subsequently processed through MLP layers and a softmax function to produce classification scores. Thus, the Particle Transformer can be viewed as a graph neural network operating on a fully connected graph, where the interactions facilitated by the attention mechanism correspond to the edge features.

### v2

@@ 4.1.2 Tokenization
The mainstream Byte Pair Encoding(BPE) [49] can introduce ambiguity and inconsistency in tokenizing numbers, leading to different segmentations of the same number based on context. For instance, the number 12345 might be tokenized as ‘12’, ‘34’, and ‘5’ in one context, or as ‘1’, ‘23’, and ‘45’ in another, losing the inherent meaning of the original numeric value. Additionally, BPE results in fragmented token IDs for numerical entities, such as ‘7’ and ‘8’ being assigned token IDs 4779 and 5014, respectively. This discontinuity in token IDs complicates the management and processing of numerical data, especially when sequential or patterned token IDs are beneficial. Similarly, single-digit tokenization [17], while straightforward, also leads to discontinuous token IDs for multi-digit numbers, such as 15 being broken down into separate tokens ‘1’ and ‘5’, each of which is then mapped to an independent token ID. This fragmentation can disrupt the continuity of numerical information, potentially making it more challenging for the model to capture the inherent structure and relationships within multi-digit numbers.
@@
  • Preservation of Numerical Meaning: By converting numbers into byte sequences, the inherent meaning of the original numeric value is preserved.
  • Consistent Tokenization: The byte tokenization method ensures that the same number is always tokenized in the same way, regardless of context, avoiding the inconsistencies seen in BPE.
  • Continuous Token IDs: Byte tokenization results in a continuous and systematic mapping of numerical values to token IDs. By employing specific computer encoding techniques, numbers like 7 and 8 can be converted into sequential token IDs, thereby streamlining the management and processing of numerical data.
Our Byte Tokenization method operates on the principle of encoding all input data into byte arrays, which serve as a common ground for further processing. The general workflow consists of four main steps:
@@
    • Text: Characters are translated into their corresponding Unicode byte values.
    • Numbers: Numbers can be processed in two distinct manners. First, a number can be treated as a string and subsequently encoded using UTF-8. For instance, the number 12345 would be represented as the string ‘12345’ and then encoded into a byte array[49, 50, 51, 52, 53]. This method is particularly useful for preserving the exact format and any leading zeros that might be significant. Second, a number can be converted into its numerical form (e.g., an integer) and then transformed into a byte array. The integer 12345 can be directly converted into a byte array representation[0, 0, 0, 0, 11, 17, 20, 21]. This approach is more efficient for arithmetic operations and when the numerical value itself is crucial.
@@
    • Image:Images are typically represented as pixel values, where each pixel can have one or more bytes depending on the color depth. For grayscale images, each pixel is represented by a single byte (8-bit), while for color images (e.g., RGB), each pixel is represented by three bytes (24-bit). The conversion process involves reading the image file and extracting the pixel values into a byte array. Consider a small 2x2 grayscale image, its pixel values, represented as $[128,64,192,255]$, would correspond to the byte array $[128,64,192,255]$ upon conversion.
@@ 4.1.3 Training
In accordance with the tokenization process outlined in Section 4.1.2, the data undergoes a series of preprocessing steps to transform jet dataset into bytes patch suitable for model training. We can present a simplified Figure 7 to illustrate how a jet is converted into a series of bytes patch. To represent jet information in a simple and effective manner, we perform a uniform formatting process on the raw data. In this process, each particle is denoted by an index number, followed by a colon that indicates the attribute values of the particle. These attribute values are arranged in the same order across different particles. These attribute values of different particles are uniformly converted into byte sequences. The byte sequences are then segmented into byte patches of a fixed length defined by the Patch Size, resulting in a total of Patch Sequence Length byte patches. Each byte patch has an input dimension equal to the patch size, and these patches form the input bytes for the model.
Data preprocessing can be implemented in either streaming or non-streaming mode, depending on the dataset size and available memory. In non-streaming mode, all data is loaded and processed at once, making this approach suitable for smaller datasets or when sufficient memory is available. Streaming mode processes data in batches as it is read from the source, which is ideal for large datasets or limited memory scenarios. This method optimizes memory usage and processing efficiency.
Training Objective The JoI task is ideally suited for classification models as it entails assigning input data to specific categories. This classification task requires the model to accurately identify and differentiate between various types of jets based on their features and attributes. Since the model need to determine what type of jet this is,the label are ‘B-jet’, ‘B-bar-jet’, ‘C-jet’, ‘C-bar-jet’, ‘S-jet’, ‘S-bar-jet’, ‘U-jet’, ‘U-bar-jet’, ‘D-jet’, ‘D-bar-jet’ and ‘G-jet’.
@@ 4.2 ParticleNet and Particle Transformer
ParticleNet [50] is based on the EdgeConv operation, which treats a jet as a graph with particles as vertices and connects each particle to its k nearest neighbors to form edges. An MLP is applied to these edges three times, updating vertex features with edge information, allowing information exchange between vertices and edges. The architecture includes three EdgeConv blocks, a global average pooling block, and two fully connected layers with a softmax function for jet predictions.
The Particle Transformer [46] uses an encoder-only design with a projection layer for classification scores. It takes particle features and interaction terms between particles as inputs, both of which are embedded using an MLP. The particle embeddings go through L attention blocks with multi-head self-attention, while interaction embeddings enhance the attention mechanism. A class token collects information from all particles via attention and is then passed through an MLP and softmax to produce classification scores.